OpenStudy (anonymous):

I have a sort of general question regarding the exponential function (ab^t). I understand what each variable stands for, but I don't understand why it works. For example: a) 100 + (100 (.10)) = 110 b) 110 + (110 (.10)) = 121 Using ab^t saves you the trouble of going through step a, above. 100 (1.10)^2 = 121 I like the fact that it works, but I don't understand why it works. Why does raising ab to the t power always work? I have the formula memorized, but I just can't see why it works. Is there a proof of this out there somewhere or can someone explain why it works? Thanks, Okubow

OpenStudy (dumbcow):

it has to do with property of exponents 2^3 = 2*2*2 i could repeat this multiplication x times or i could say 2^x from above example 100(1.1)^2 = 100(1.1*1.1) this is same as doing 100*1.1 then multiplying by 1.1 again Or multiply 1.1*1.1 first then multiply by 100 The key is i get the same answer no matter which order i do the multiplication using exponents is just a short hand way of doing repeated multiplications im sure there is some fancy mathematical proof out there anyway thats the gist of it

OpenStudy (anonymous):

Thanks, dumbcow. That's the kind of answer I was looking for. This helped me to visualize it!: 100(1.1)^2 = 100(1.1*1.1)

OpenStudy (dumbcow):

good to hear your welcome